4 times the measure of the complement angle is equal to 2/3 the measure of the original angle's supplement. What is the measure of the angle?
Answer by prateekagrawal(56)

let the angle be x.
then
complement of that = 90-x
so, 4 times of complement = 4(90-x)
2/3 of original angle's supplement= 2/3 (180-x)
so,
4(90-x)= 2/3 (180-x)
cross multiply:
12(90-x) = 360 - 2x
12*90 - 12x = 360 - 2x
1080- 360= 12x-2x
720 = 10x
x= 72.
is the answer.